(a) Find the uncertainty in the position of an electron in an infinite square-well potential if the electron is in the n=5 state and the box is 0.10nm wide.
(b) Find the uncertainty in the momentum of an electron in an infinite square-well potential if the electron is in the n=5 state and the box is 0.10nm wide.
Answer #1

well poten hae men b we asug ,wen, n= 5 and L=o.lonm o.Ixlown 12 (5) 240-H 2960.89 I S 一13 O00 913X lo - 18-4 -22nase une n=S -o.lonm =0./ox10 vn / 古:h -34